# Comparing Specialised Small and General Large Language Models on Text Classification: 100 Labelled Samples to Achieve Break-Even Performance

### v2

@@ 3 Investigating the Impact of Dataset Size
Figure 2: The impact of varying size of available labelled training samples (in logarithmic scale) on the performance of fine-tuning, prompting, in-context learning and instruction-tuning approaches across the binary (SST2, MRPC, CoLA, BoolQ) and multi-class datasets (AG News, TREC, SNIPS, DB Pedia), reported using F1 macro and its standard deviation. Although the impact is dataset dependent (with binary datasets requiring more labelled samples), we can observe that fine-tuning can outperform large language models used in zero/few-shot setting when trained only small number of labelled samples ($10-1000$). On the other hand, instruction-tuning of small (Flan-T5) or medium (Mistral/Zephyr) language models with only 10 labelled samples and using them in zero/few-shot setting often leads to performance on par, or better, than the one in much larger language models (LLaMA-2/ChatGPT) used in zero/few-shot setting without additional tuning.
3.1 Comparison of Specialised and General Models
In this section, our goal is to answer the following main research question: RQ1: How do the specialised models compare to general models on average as the number of labelled samples increases? To answer this research question, we identify the first break-even points (average performance) in the performance of specialised (fine-tuning, instruction-tuning) models in comparison to general ones (prompting, in-context learning), but also with each other. These break-points specify the point after which the performance of fine-tuning is better on average, but may still be lower due to performance variance, such as when the randomness is not sufficiently addressed (low number of runs is used or runs are cherry-picked). The aggregated results are presented in Figure 1, with the results for individual models and datasets presented in Figure 2.
Specialised models can outperform the general ones using only small number of labelled examples. To outperform zero-shot setting (prompting), fine-tuning approaches require on average between $10-1000$ labelled samples. In case of few-shot setting (in-context learning), the number of required labelled samples is higher, on average between $100-5000$ samples.
The instruction-tuned models (small or medium sized) represent a good balance between the generality of the models and the samples required for specialisation. To outperform these models, fine-tuning approaches often require a majority of the labelled dataset. In addition, we observe a consistent performance of instruction-tuned models, regardless of how many labelled samples are used. The performance achieved by instruction-tuned models with $10$ labelled samples is close to the one achieved with the full labelled dataset. Furthermore, they consistently outperform all the general larger counterparts. As such, instruction-tuned models achieve the almost-best performance with only a fraction of labelled samples required by fine-tuning, making them an ideal specialised models in many cases.
However, the number of required labelled samples and the comparison between models is heavily dependent on the dataset and task characteristics, such as how many classes are used (binary vs. multi-class setting), length of sentences (e.g., SST2 vs. CoLA), whether the task requires working with a single sentence or a paired/multiple inputs (e.g., SST2/CoLA vs. MRPC/BoolQ), or the overall type of the task (e.g., simple sentiment classification vs. question answering) that is defined by these characteristics. The largest difference can be observed between binary and multi-class datasets. On multi-class datasets, fine-tuning requires only up to $100$ labelled samples to outperform any of the models (even instruction-tuned ones). On binary datasets the required samples are as high as $6000$ to outperform zero and few-shot setting, or up to $16000$ to outperform the instruction-tuned models. The remaining characteristics have lower, less consistent impact. For example, the task type on binary datasets significantly affects the samples, with BERT on BoolQ not being able to outperform majority of the remaining approaches even with full dataset, but has almost no effect on multi-class datasets. At the same time, some datasets cause significant problems to specific models. For example instruction-tuning of Flan-T5 on CoLA dataset, prompting of Mistral-7B and LLaMA-2 models on TREC dataset, Zephyr-7B on BoolQ dataset, or Mistral-7B on MRPC dataset, or in-context learning of Flan-T5 model on DB-Pedia dataset.
Similarly to instruction-tuning, general models show consistent performance regardless of the number of labelled samples. Only in-context learning shows slight fluctuations in performance. However, this fluctuation is caused only indirectly by the number of labelled samples, as it affects the choice of in-context examples (i.e., what samples are available). As such, only the performance of fine-tuning consistently improves as more labelled samples are used. In addition, we do not observe consistent improvement of in-context learning (few-shot setting) over prompting (zero-shot setting). Instead, the performance difference is more dataset dependent. On the multi-class datasets, in-context learning is beneficial over prompting in majority of the cases (only exception is Flan-T5 on DB-Pedia and ChatGPT on AG News). On the binary datasets, the benefit of in-context learning strongly depends on the models. For example on BoolQ dataset, Zephyr-7B benefits from in-context learning while Mistral-7B does not. At the same time, the improvement of in-context learning over prompting is minor in some cases (such as the increase in Flan-T5 on the SNIPS dataset from $42.5\%$ to $44.5\%$).
Table 2: Break-even points between fine-tuning and the remaining approaches on all investigated datasets. The required labelled samples are indicated only for the best performing model from the 3 groups of models of different sizes: 1) Flan-T5 (small); 2) Mistral/Zephyr (medium); and 3) LLaMA2/ChatGPT (large). The first break-even point (average performance) and the second one (average and variance) are separated with the "/" symbol, with superscript indicating the percentual increase and "NA" indicating no break-even point exists (fine-tuning never outperforms the other approach). We observe a significant influence of the variance in results on the number of required labelled samples, leading to an increase of up to 1500% (from $500$ to $8000$ samples).
Prompting In-Context Learning Instruction-Tuning
Flan-T5 Mistral/Zephyr LLaMA2/ChatGPT Flan-T5 Mistral/Zephyr LLaMA2/ChatGPT Flan-T5 Mistral/Zephyr
SST2 BERT $90/125^{+39\%}$ $25/50^{+100\%}$ $500/1200^{+140\%}$ $125/225^{+~{}80\%}$ $75/150^{+100\%}$ $3500/10000^{+186\%}$ $16000/20000^{+25\%}$ $4000/20000^{+400\%}$
RoBERTa $80/100^{+25\%}$ $50/75^{+~{}50\%}$ $300/1000^{+233\%}$ $100/225^{+125\%}$ $75/175^{+133\%}$ $5500/~{}~{}8000^{+~{}45\%}$ $13000/20000^{+54\%}$ $2500/20000^{+700\%}$
MRPC BERT $600/900^{+50\%}$ $450/700^{+55\%}$ $900/1750^{+94\%}$ $700/1100^{+57\%}$ $600/1250^{+108\%}$ $650/1250^{+92\%}$ $900/1750^{+~{}94\%}$ $1500/4000^{+167\%}$
RoBERTa $225/350^{+55\%}$ $175/300^{+71\%}$ $300/~{}~{}550^{+83\%}$ $250/~{}~{}400^{+60\%}$ $225/~{}~{}400^{+~{}78\%}$ $225/~{}~{}400^{+78\%}$ $300/~{}~{}600^{+100\%}$ ${}~{}400/1000^{+150\%}$
CoLA BERT $10/10^{+~{}~{}~{}0\%}$ $20/~{}~{}75^{+275\%}$ $400/800^{+100\%}$ $50/100^{+100\%}$ $850/NA$ $350/2000^{+471\%}$ $10/10^{+~{}~{}0\%}$ $3000/NA$
RoBERTa $10/25^{+150\%}$ $75/100^{+~{}33\%}$ $250/750^{+200\%}$ $75/150^{+100\%}$ $800/NA$ $250/1500^{+500\%}$ $10/15^{+50\%}$ $1750/NA$
BoolQ BERT $NA/NA$ $1800/3500^{+94\%}$ $NA/NA$ $NA/NA$ ${}~{}500/8000^{+1500\%}$ $NA/NA$ $NA/NA$ $NA/NA$
RoBERTa $4250/6000^{+41\%}$ ${}~{}800/1250^{+56\%}$ $4750/5000^{+5\%}$ $1750/4000^{+129\%}$ $1000/3000^{+~{}200\%}$ $6000/NA$ $4750/NA$ $3000/4000^{+33\%}$
AGNews BERT $50/60^{+20\%}$ $30/50^{+67\%}$ $50/60^{+20\%}$ $550/1100^{+100\%}$ $75/100^{+33\%}$ $50/100^{+100\%}$ $550/1100^{+100\%}$ $100/300^{+200\%}$
RoBERTa $40/50^{+25\%}$ $40/50^{+25\%}$ $50/50^{+~{}~{}0\%}$ $350/1100^{+214\%}$ $50/~{}~{}50^{+~{}~{}0\%}$ $50/~{}~{}50^{+~{}~{}~{}0\%}$ $350/1100^{+214\%}$ ${}~{}75/250^{+233\%}$
TREC BERT $60/80^{+33\%}$ $30/40^{+33\%}$ $50/75^{+50\%}$ $90/125^{+39\%}$ $75/100^{+33\%}$ $60/90^{+50\%}$ $100/150^{+50\%}$ $100/175^{+75\%}$
RoBERTa $40/50^{+25\%}$ $30/35^{+17\%}$ $40/50^{+25\%}$ $55/~{}~{}75^{+36\%}$ $50/~{}~{}60^{+20\%}$ $40/60^{+50\%}$ ${}~{}60/~{}~{}90^{+50\%}$ ${}~{}75/100^{+33\%}$
SNIPS BERT $25/40^{+60\%}$ $10/20^{+100\%}$ $40/50^{+25\%}$ $25/40^{+~{}~{}60\%}$ $50/75^{+50\%}$ $50/70^{+40\%}$ $25/40^{+~{}~{}60\%}$ $75/175^{+133\%}$
RoBERTa $25/40^{+60\%}$ $10/30^{+200\%}$ $30/40^{+33\%}$ $20/40^{+100\%}$ $40/50^{+25\%}$ $35/50^{+43\%}$ $20/40^{+100\%}$ $50/~{}~{}60^{+~{}20\%}$
DBPedia BERT $50/70^{+40\%}$ $40/50^{+25\%}$ $70/80^{+14\%}$ $10/10^{+0\%}$ $100/200^{+100\%}$ $80/200^{+150\%}$ $125/225^{+~{}~{}80\%}$ $150/250^{+~{}~{}67\%}$
RoBERTa $45/45^{+~{}~{}0\%}$ $40/45^{+13\%}$ $50/50^{+~{}~{}0\%}$ $10/10^{+0\%}$ ${}~{}70/100^{+~{}~{}43\%}$ $50/100^{+100\%}$ ${}~{}80/160^{+100\%}$ $100/225^{+125\%}$
3.2 Impact of Performance Variance on Comparison
In this section, we answer following research question: RQ2: How does the variance from repeated runs affect the comparison between specialised and general models? Our goal is to determine how the number of required samples increases when the performance variance is taken into consideration. To accomplish this, we identify the second break-even points (average+variance) denoting "worst-case sizes" of training sets. They represent the point after which fine-tuning shows better performance, even when the variance is taken into consideration (i.e., there is low probability that the performance would be worse when using low number of runs). The results from this investigation are presented in Figure 2 (aggregated in Figure 1) and the comparison of first and second break-even points for the best performing models of different sizes in Table 2.
Performance variance has a significant effect on the break-even points between models, increasing the number of required labelled samples by a large amount. On average, fine-tuning requires $2-3$ times more labelled samples (increase of $100-200\%$) to outperform the remaining approaches when taking the variance into consideration. In specific cases, the impact of variance is significantly lower (negligible), as we observe a $0\%$ increase in the number of required samples (e.g., RoBERTa and LLaMA2/ChatGPT prompting and in-context learning on AG News dataset). At the same time, the impact of variance is significantly higher in other cases, leading to an increase of up to $1500\%$, or an increase where even with the full labelled dataset the second break-even point is not achieved (e.g., BERT and RoBERTa in comparison to Mistral/Zephyr in-context learning on the CoLA dataset).
However, the impact of the performance variance strongly depends on the dataset and the overall number of labelled samples required for the first break-even point. Looking at the increase in absolute numbers instead of a percentual increase, an increase of $100-200\%$ represents the need to annotate only between $10-750$ more samples in some cases (e.g., most multi-class datasets, such as BERT or RoBERTa in comparison to Flan-T5 in-context learning on AG News dataset), while representing an increase of $2000-6500$ or more labelled samples in other cases (e.g., binary datasets, such as RoBERTa in comparison to Mistral/Zephyr in-context learning on BoolQ dataset, or BERT in comparison to LLaMA2/ChatGPT in-context learning on SST2 dataset).
In addition, the second break-even point heavily depends on the variance of the approaches/models used. In case of prompting, which shows the lowest performance variance, the highest observed increase is $233\%$ (RoBERTa with LLaMA2/ChatGPT in-context learning on SST2 dataset) in relative numbers or $850$ (BERT with LLaMA2/ChatGPT on MRPC dataset) in absolute numbers. On the other hand, the increase in the number of required labelled samples is significantly higher for in-context learning approaches (which show the highest variance across the runs) or specific cases in the instruction-tuning approaches (where the variance is more dependent on the model, with larger models showing higher variance), with the highest observed increase of $1500\%$ (BERT with Mistral/Zephyr in-context learning on BoolQ dataset) in relative numbers or increase of $17500$ labelled samples (RoBERTa with Mistral/Zephyr instruction-tuning on SST2 dataset) in absolute number. The variance in the fine-tuning approaches does not have as much of an effect, as it is strongly affected by the number of labelled samples – the variance is higher with low number of labelled samples and almost non-existent (lower than prompting) with majority of the dataset (more than $\sim 60\%$ of the labelled samples in the dataset), with the exception of binary dataset where even with all the labelled samples, the variance is quite high. In some cases, instruction-tuning leads to lowered variance of the model (in comparison to using it in zero/few shot setting), but in other cases leads to even larger variance in results by introducing additional sources of randomness in the optimisation process.
3.3 Additional Insights from the Investigation
In this section, we are mainly interested in how other factors such as the size of the model affect the comparison (both in terms of overall performance and variance).
@@ 3.3 Additional Insights from the Investigation
Larger models do not consistently lead to better performance across different approaches. With the fine-tuning, we observe the expected behaviour. On lower number of labelled samples (up to $20-100$ depending on the dataset) the smaller BERT model is able to outperform the larger RoBERTa model. However, the larger model can achieve the highest performance on the different datasets using fewer labelled samples. As such, the individual break-even points appear earlier for the larger fine-tuning models. On the other hand, in case of prompting, in-context learning and instruction-tuning, the smaller models are often able to achieve better or similar performance than the larger counterparts even though the different in the number of parameters is large. For example, Flan-T5 in-context learning on the MRPC and CoLA dataset outperforms in-context learning with LLaMA2 and ChatGPT models, while on the BoolQ dataset the situation is opposite. Similarly, instruction-tuning Mistral-7B and Zephyr-7B models often leads to similar or lower performance than the instruction-tuning of Flan-T5 model. This behaviour does not necessarily depend on the dataset characteristics. On some datasets that can be considered harder (longer inputs, more classes, harder tasks that require better language understanding), the Flan-T5 model in zero/few shot setting still outperforms the larger models, while on other datasets it underperforms them. The only impact of the dataset characteristics is the probability of a ’failure mode’, where the smaller model does not perform at all at the task. For example, using in-context learning with Flan-T5 on the DB-Pedia dataset always leads to prediction of a single class for all of the samples, with this behaviour disappearing when reducing the number of classes, or using prompting. Similarly, in-context learning with Flan-T5 on the BoolQ dataset (which is characteristic with its longer input) shows a significantly higher variance. We believe the main culprit for this is the limited context size of the smaller model.
@@
As such, the number of required labelled samples is not consistently dependent on the size of the general models. For prompting and in-context learning, we observe many cases when the break-even point for the smaller Flan-T5 requires similar or larger number of labelled samples than the larger models. For example, on the AG News dataset the number of required labelled samples for RoBERTa to outperform Flan-T5 in-context learning is $550$ (or $1100$ when taking the variance into consideration), while for the medium models it is $75$ (or $100$) and for large models it is $50$ (or $100$). Similar observations are present in instruction tuning as well, where the BERT model requires $16000$ samples to outperform Flan-T5 instruction-tuning, but only $4000$ to outperform medium sized models. Curiously, fine-tuning can often outperform the medium sized models (Mistral/Zephyr) with lower number of samples than the smallest Flan-T5 model. In addition, on the binary datasets, fine-tuning consistently requires the highest number of samples to outperform the largest models (LLaMA-2 and ChatGPT).
@@
In addition, using larger models does not necessarily lead to lower performance variance. With the fine-tuning models, we observe the expected behaviour again. Larger model shows lower variance when using enough labelled data, but shows larger variance than the smaller model when dealing with limited data. The performance variance in prompting is similar across all the models, as there are almost no sources of randomness that could affect it. On the other hand, the variance in in-context learning is not as consistent. Comparing the smallest model (Flan-T5) and the largest models (LLaMA-2 and ChatGPT), we observe similar performance variance on majority of the datasets. Only on specific datasets the variance of Flan-T5 is significantly higher (e.g., BoolQ) or significantly lower (AG News). At the same time, the variance of the medium sized models (Mistral-7B and Zephyr-7B) is often significantly higher than their smaller or larger counterparts. Finally, with instruction-tuning, we observe that the Mistral-7B and Zephyr-7B models often show larger variance when compared to Flan-T5. Although this may be due to the different type of instruction-tuning (full tuning in Flan-T5 vs. LoRA tuning in Mistral/Zephyr), we observe similar behaviour of these models in the zero/few shot setting as well.
3.4 Recommendations Based on the Findings
Based on the findings, we argue that using general large language models in a zero or few shot setting is a preferable solution only in specific cases. First, if the task does not have well defined classes (e.g., planning or answering free text questions) or the task requires some kind of generation (e.g., translation, summarisation). Even though the instruction-tuning would provide benefit for such setting, the general models are often already tuned for this behaviour and further tuning would require extensive dataset. As providing in-context examples to allow for in-context learning could be problematic in such a case, prompting is more efficient option. Second, if faced with a limited annotation and computation budget, where we either cannot obtain large enough labelled dataset for fine-tuning or we are not able to instruction-tune the general models. Third, if we are interested in quick prototyping and approximation of the overall performance we can achieve on the task, before deciding whether to dedicate more time and budget for specialising the models. This recommendation is based on our main finding (specialised models require only small number of labelled samples to outperform general models) and is in contrast to the common practice of evaluating prompting and in-context learning on classification tasks Liu et al. (2023).
Fine-tuning of smaller models is preferable if we have large annotation budget, but small computation budget (and a classification task), or if the task is not well designed for generation and would require additional modifications (e.g., complicated mapping of the generated text to classes, or when multiple classes are possible for each sample). Even when fine-tuning small model (BERT), such approach can outperform all the general models if the annotation budget is large enough. In this case, training the largest possible model (allowed by the computation budget) should be opted for, to reduce the number of required labelled samples as much as possible, to achieve the highest possible performance and to prevent any shortcomings on tasks that require better language understanding. Based on our findings, we need to annotate only approximately $1000$ samples, with larger models requiring fewer samples.
Instruction-tuning of the larger language models is the optimal solution as it provides consistent benefits for all the other cases (large computation budget with either small or large annotation budget). Even with small number of labelled samples ($10$), instruction-tuning slightly larger general language model (i.e., models that were already pre-trained for prompting and in-context learning such as Flan-T5) can lead to the best overall performance. Instruction tuning provides the best trade-off for the performance and the required resources, as increasing the number of labelled samples does not have a significant impact on the performance, while increasing the size of the instruction-tuned model has only low impact on performance. However, if interested in best overall performance on the task, using the largest general model and instruction-tuning it with large number of labelled samples should provide the most benefit even at the cost of significant increase in effort (and strong diminishing returns).
Finally, when comparing between different approaches and models, the performance variance should be taken into consideration, as it has a strong impact on the comparison. If only a single run (or low number of runs) is used, one of the models can be incorrectly denoted as better only based on a random chance. The use of multiple runs is especially important when using in-context learning, instruction-tuning or fine-tuning on low number of labelled samples.
4 Conclusion
In this paper, we investigate the impact of changing the number of available labelled training samples on the performance and comparison of different, currently popular, approaches for data-efficient learning in NLP. The main focus of the investigation is to determine how many labelled training samples are needed for the specialised small language models (obtained through fine-tuning or instruction-tuning) to outperform their general larger counterparts used in zero and few-shot settings. Based on the break-even points from 8 representative text classification datasets of various characteristics, we find that the number of required training samples is quite small, but strongly dependent on the dataset and task characteristics, with more labelled samples needed for binary datasets and tasks that require better language understanding. In addition, we identify a significant influence of the performance variance, stemming from the sensitivity of the different approaches especially in-context learning and fine-tuning with few labelled samples, on the overall comparison and the number of required labelled samples. We can conclude that the specialised smaller models are a strong contender on the text classification tasks, with the general language models showing their benefits for specific cases, such as quick prototyping or when faced with extremely limited annotation or computation budget. Finally, based on findings of our experiments, we provide recommendations that allow for better decision making regarding what approach to use for different settings.
Limitations
The investigation is done on a set of English classification datasets with various characteristics (number of classes, input size, task type, etc.). This choice may limit the generalisation of our findings to other datasets, tasks and languages, such as the generation tasks, which are more representative for the general models and on which the fine-tuning cannot be used as easily (e.g., question answering, summarisation, translation). However, we explicitly discuss this in Section 3.4.
In addition, we perform the investigation across a lower number of smaller models. We work with the base versions of the smaller BERT, RoBERTa and Flan-T5 models. The larger models used, specifically Mistral-7B, Zephyr-7B and LLaMA-2 13B, are all used with the 4 bit quantisation. The largest model we use is the ChatGPT (3.5 version). We explicitly choose these models to provide a more contrasted comparison between the performance while limiting the required computation costs and to provide more in-depth analysis of their behaviour (including how the performance variance affects the comparisons). Even though the fine-tuning models consistently outperforms the larger counterparts with low number of labelled samples, the choice of smaller models may limit the generalisation of our findings. However, we publicly release the source code of our investigation to allow for its extension to other models and datasets.
Due to the significant computation costs incurred by the inference of larger language models (namely LLaMA-2, Mistral-7B and Zephyr-7B), instruction-tuning of the medium sized models (Mistral/Zephyr) and the additional other costs of the ChatGPT model, we evaluate the models in a limited setting. The evaluation is done on a randomly selected set of 1 000 samples (following the practice in prior works Gao et al. (2021); Chang and Jia (2023); Sclar et al. (2024); Li and Qiu (2023); Köksal et al. (2023)). As the majority of the datasets is smaller (e.g., MRPC, BoolQ, CoLA, SNIPS, TREC), this decision may not be as problematic, as the used 60:20:20 split leads to approximately the 1000 test samples. However, on the larger datasets (SST2, AG News and DB Pedia) this decision may skew the results. In addition, we run the ChatGPT prompting and in-context learning only 10 times instead of 20 and only for the 1 000 labelled training samples setting, due to its costs.
Finally, we use the same prompt for all the models, which is a result of a prompt engineering on the Mistral-7B model. The prompt was created based on dataset description, the prompts used in related works and the formats recommended for different models (e.g., taking inspiration from Sun et al. (2023)). As such, this may not represent the optimal format for all the models (as identified in previous works Sclar et al. (2023)) and performing the investigation on multiple different prompts may improve the overall model performance and affect the findings. However, we opted for using only a single optimised prompt format to reduce the computation costs and provide more in-depth analysis on larger number of models.
Ethical Considerations
The experiments in this paper work with publicly available benchmark datasets GLUE and SuperGLUE and other publicly available datasets (AG News, TREC, SNIPS and DB Pedia), citing the original authors. As we were not able to determine the license for the tasks used, we opted to use them in as limited form as possible, adhering to the terms of use (no annotation of the test set) defined by the GLUE and SuperGLUE and applying it to other datasets as well. We do not work with any personally identifiable information or offensive content and perform no crowdsourcing for further data annotation. In addition, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the advancement of the field of machine learning. Finally, we follow the license terms for all the models we use (such as the one required for the use of the LLaMA-2 model). It is possible the large language models we use contain biases and potentially offensive or harmful content. However, the original authors of these models reduce this bias as much as possible. At the same time, we do not release any output of the models which should further reduce the potential bias and negative impact.
Impact Statement: CO2 Emissions Related to Experiments

### v3

@@ Appendix B Effect of Model Size on Comparison and Dataset Dependence
Figure 4: The impact of varying size of available labelled training samples (in logarithmic scale) on the performance of fine-tuning, prompting, in-context learning and instruction-tuning approaches across the binary (SST2, MRPC, CoLA, BoolQ) and multi-class datasets (AG News, TREC, SNIPS, DB Pedia), reported using F1 macro and its standard deviation. For each approach, we group the models based on their size into small (Flan-T5), medium (Mistral/Zephyr) and large (LLaMA/GPT) and report the best performing model for each group. Even though the effect of model size is often significant, it does not follow common assumption. The smaller models, especially in prompting, in-context learning and instruction-tuning, often achieve better performance than the medium or large general models.
Figure 5: The impact of varying size of available labelled training samples (in logarithmic scale) on the performance of fine-tuning, prompting, in-context learning and instruction-tuning approaches, reported using F1 macro and its standard deviation. All the models are presented for each approach.
Previous studies have observed that the size of the model significantly affects its ability to perform prompting and in-context learning. As such, in this section, we are interested in how the size of the model and other factors affect the comparisons, both in term of overall performance and variance, but also other properties as well. Instead of reporting results only for the best performing model for each approach, we group the models based on their size into small (Flan-T5), medium (Mistral/Zephyr) and large (LLaMA/GPT) and report the best performing model for each group. The results are included in Figure 4. In addition, we provide the results for the individual models for each approach and dataset in Figure 5.
@@
First of all, larger models do not consistently lead to better performance across different approaches. With the fine-tuning, we observe the expected behaviour. On lower number of labelled samples (up to $20-100$ depending on the dataset) the smaller BERT model is able to outperform the larger RoBERTa model. However, the larger model can achieve the highest performance on the different datasets using fewer labelled samples. As such, the individual break-even points appear earlier for the larger fine-tuning models. On the other hand, in case of prompting, in-context learning and instruction-tuning, the smaller models are often able to achieve better or similar performance than the larger counterparts even though the different in the number of parameters is large. For example, in-context learning with small models (Flan-T5) on the AG News and SST2 dataset outperforms in-context learning with medium (Mistral-7B and Zephyr-7B) or large (LLaMA and GPT) models, while on the BoolQ dataset the situation is opposite. Similarly, instruction-tuning with medium models (Mistral-7B and Zephyr-7B) models often leads to similar or lower performance than the instruction-tuning of the small (Flan-T5) model.
This behaviour does not necessarily depend on the dataset characteristics. On some datasets that can be considered harder (longer inputs, more classes, harder tasks that require better language understanding), the Flan-T5 model in zero/few shot setting still outperforms the larger models, while on other datasets it underperforms them. The main impact of dataset characteristics is causing a ’failure mode’ for specific models, where the smaller model does not perform at all at the task. For example, using in-context learning with small (Flan-T5) model on the DB-Pedia dataset always leads to prediction of a single class for all of the samples, with this behaviour disappearing when reducing the number of classes, or using prompting. Similarly, in-context learning with the smaller (Flan-T5) model on the BoolQ dataset (which is characteristic with its longer input) shows a significantly higher variance. We believe the main culprit for this is the limited context size of the smaller model.
The impact of the model size is the most explicit for fine-tuning. On specific datasets that can be considered harder, the smaller BERT model performs significantly worse. For example, the BERT model on the BoolQ fails to outperform majority of the remaining approaches even when using the full dataset. However, this impact is minimal on the multi-class datasets.
@@
Overall the number of required labelled samples is not consistently dependent on the size of the general models. For prompting and in-context learning, we observe many cases when the break-even point for the smaller model requires similar or even larger number of labelled samples than the medium or large models. For example, on the AG News dataset the number of required labelled samples for RoBERTa to outperform Flan-T5 in-context learning is $300$ (or $800$ when taking the variance into consideration), while for the medium models it is only $100$ and for large models it is $50$. Similar observations are present in instruction tuning as well. Curiously, fine-tuning can often outperform the medium sized models (Mistral/Zephyr) with lower number of samples than the smallest Flan-T5 model. In addition, on the binary datasets, fine-tuning consistently requires the highest number of samples to outperform the largest models (LLaMA and GPT).
The benefit of increasing the number of in-context examples for in-context learning is affected by the size of the models as well. However, this benefit is, again, dataset dependent and is closely tied to the context length and the dataset characteristics. For example, on the binary datasets with shorter sentences (SST2/CoLA), the benefit of more samples is significantly higher than on the datasets with multiple classes or with longer sentences (BoolQ/DB Pedia). This also directly affects the benefit of instruction-tuning. Although the benefit is larger for the medium and large models, in many cases increasing the number of in-context samples only leads to these models to achieve performance similar to the one achieved by the smaller model – i.e., for in-context learning with medium models, we require up to $100$ samples to achieve performance comparable to small models that use only 10 labelled samples.
In addition, we do not observe consistent improvement of in-context learning (few-shot setting) over prompting (zero-shot setting). Instead, the performance difference is more dataset dependent. On the multi-class datasets, in-context learning is beneficial over prompting in majority of the cases. On the binary datasets, the benefit of in-context learning strongly depends on the models. At the same time, the improvement of in-context learning over prompting is minor in some cases (such as the increase in Flan-T5 on the SNIPS dataset from $42.5\%$ to $44.5\%$).
@@
Finally, using larger models does not necessarily lead to lower performance variance. With the fine-tuning models, we observe the expected behaviour again. Larger model shows lower variance when using enough labelled data, but shows larger variance than the smaller model when dealing with limited data. The performance variance in prompting is similar across all the models, as there are almost no sources of randomness that could affect it. On the other hand, the variance in in-context learning is not as consistent. Comparing the smallest model (Flan-T5) and the largest models (LLaMA and GPT), we observe similar performance variance on majority of the datasets. Only on specific datasets the variance of Flan-T5 is significantly higher (e.g., BoolQ) or significantly lower (AG News). At the same time, the variance of the medium sized models (Mistral-7B and Zephyr-7B) is often significantly higher than their smaller or larger counterparts. Finally, with instruction-tuning, we observe that the medium models (Mistral-7B and Zephyr-7B) often show larger variance when compared to Flan-T5. Although this may be due to the different type of instruction-tuning (full tuning in Flan-T5 vs. LoRA tuning in Mistral/Zephyr), we observe similar behaviour of these models in the zero/few shot setting as well.
Appendix C Effect of Quantisation on General Models
Figure 6: The comparison between 4-bit quantised and non-quantised Mistral-7B and Zephyr-7B models used for prompting across all datasets. Even though the non-quantised models achieve better performance, the difference in performance is minimal in almost all cases. In addition, the impact on the variance is negligible.
Figure 7: The comparison between 4-bit quantised and non-quantised Mistral-7B and Zephyr-7B models used for in-context learning across all datasets. The impact of quantisation is not consistent across datasets. The quantised models often achieve better performance than non-quantised ones, with the difference being often small. In addition, the impact on the variance is negligible.
As the general large language models contain high number of parameters, it is a common practice to use their quantised versions instead to reduce the computation costs. Although some paper research focuses on observing how this affects the performance, the impact of quantisation on the comparison between models and the sensitivity to the effects of randomness is not well understood. Therefore, in this section, our aim is to determine how the use of quantised models affect the comparison and the overall sensitivity to the effects of randomness. To achieve this, we run the same experiment with 4-bit quantised and non-quantised versions of the Mistral-7B and Zephyr-7B models using the prompting and in-context learning and compare their average performance and variance. The results of this comparison are presented in Figure 6 for prompting and in Figure 7 for in-context learning.
The non-quantised models perform slightly better than 4-bit quantised models when it comes to prompting. The difference in performance between the 4-bit quantised and non-quantised models is minimal in almost all the cases, with non-quantised models achieving slightly higher performance. Only exception is the MRPC dataset, where the Zephyr 4-bit quantised model achieves better performance, and DB Pedia where both Mistral and Zephyr achieve higher performance. On the other hand, the non-quantised version of the Zephyr model achieves significantly higher performance on the BoolQ, CoLA and AG News datasets.
The impact of quantisation is less consistent for in-context learning. In contrast to prompting, the 4-bit quantised models used for in-context learning outperform their non-quantised versions more often. In addition, the difference in performance in such cases is often significantly higher. At the same time, when the quantised models do not outperform the non-quantised versions, the difference in performance is often small (only exception is the Zephyr model on the CoLA dataset).
Furthermore, the impact of quantisation on the sensitivity to the effects of randomness and the performance variance it causes is negligible. In almost all cases, the observed standard deviation is similar regardless of the model version. In case the performance variance is different, the non-quantised models achieve higher performance. For example, the prompting of Zephyr model on the MRPC, BoolQ or AG News datasets. In case of in-context learning, the difference in performance variance is even less significant.
Finally, quantisation has no impact on how much the models benefit from increasing the number of in-context examples or their context length. In all cases, we observe similar increase in performance (and decrease in standard deviation) when increasing the number of examples. At the same time, we observe that both quantised and non-quantised versions of the models achieve the ‘failure state‘ at the same number of samples. We believe this failure state corresponds to reaching the context size of the models, as on the datasets with longer sentences (BoolQ/DB Pedia), it is encountered sooner than on datasets with shorter sentences (SST2/CoLA). However, in many cases, the best performance of the model is achieved well before using the whole context size of the models.
We can conclude that the use of quantised models has minimal impact on the comparison between different approaches. As such, considering compute-efficiency and performance trade-off in the comparison, the quantised models should be preferred for the general large language models whenever possible.
Appendix D Impact Statement: CO2 Emissions Related to Experiments